A man uses a rock as a fulcrum and lifts a 300-kg boulder with a metal bar 6.00 m long. Calculare the force necessary to lift the boulder if the fulcrum is:


(a) 1.00 m from where the bar contacts the rock.


(b) 3.00 m from where the bar contacts the rock

1
Expert's answer
2022-03-13T18:46:32-0400

Answer


At the equilibrium the total torque around the pivot point (fulcrum) is zero, so

"F_1 d_1=F_2 d_2"

or

"Fd_1=mgd_2"

Thus

"F=mg d_2\/d_1"

"d_2=1 m, \\d_1=l-d_2=6-1=5 m\\

F=300×9.8×1/5=588 N

\\d_2=3 m, \\d_1=l-d_2=6-3=3 m

\\F=300×9.8×3/3=2940 N"